In a large hospital an average of 3 out of every 5 patients ask for water with their meal. A random sample of 10 customers is selected. Assuming a binomial distribution, find the probability that i. Exactly 6 patients ask for water with their meal ii. At most 4 patients ask for water with their meal iii. At least 3 patients ask for water with their meal iv. Find the mean of this distribution v. Find the standard deviation of this distribution.
In a mainframe computer centre, execution time of programs follows an exponential distribution. The average execution time of the programs is 5 minutes. Find the probability that the execution time of programs is: i. Less than 4 minutes ii. More than 6 minutes
1
Expert's answer
2020-10-01T13:55:36-0400

Let "X=" number of patients who asked for water: "X\sim Bin(n, p)"

Given "p=3\/5=0.6, n=10"

i. "P(X=6)=\dbinom{10}{6}(0.6)^6(1-0.6)^{10-6}=0.250822656"


ii. "P(X\leq4)=P(X=0)+P(X=1)+P(X=2)+"

"+P(X=3)+P(X=4)="

"=0.0001048576+0.001572864+0.010616832+"

"+0.042467328+0.111476736=0.1662386176"


iii. "P(X\geq4)=1-P(X=0)-P(X=1)-"

"-P(X=2)-P(X=3)="

"=1-0.0001048576-0.001572864-0.010616832-"

"-0.042467328=0.9452381184"


iv. "E(X)=np=10(0.6)=6"


v. "V(X)=np(1-p)=10(0.6)(1-0.6)=2.4"

"\sigma_X=\sqrt{2.4}\approx1.55"


Let "X=" the execution time of programs: "X\sim Exponential(\lambda)"

Given "\lambda=5"

i. "P(X<4)=1-e^{-5(4)}\approx0.999999998\approx1"


ii. "P(X>6)=e^{-5(6)}\approx10^{-13}\approx0"